Angle two is 3 times angle one.
Angle three is 6 more than two times angle one

Answer by checkley77(12844)

If this is a triangle then:
b=3a
c=2a+6
a+b+c=180
a+3a+(2a+6)=180
6a=180-6
6a=174
a=174/6
a=29 degrees. for angle 1.
b=3*29=87 degrees is angle 2.
c=2*29+6=58+6=64 degrees ia angle 3. Proof:
29+87+64=180
180=180